the angle of elevation to the top of a building from a position on the ground 100 meters from the base of the building is 40. how tall is the building, to the nearest meter?
Answer by asuar010(338)

by drawing a triangle to sove this problem we deduct that the correct trigonometric functio to use tan 40= x/100 so we obtain x=100*tan(40) so the building is 73 meters (72.6 rounded up)